The stopcock connecting a 2.95 L bulb containing nitrogen gas at a pressure of 2.75 atm, and a 4.87 L bulb containing carbon dioxide gas at a pressure of 2.15 atm, is opened and the gases are allowed to mix. Assuming that the temperature remains constant, the final pressure in the system is
1
Expert's answer
2018-03-23T08:48:14-0400
The final volume: 2.95 + 4.87 = 7.82 L
P1V1 = P2V2
P2 (N2) = 2.75 atm x 2.95 L / 7.82 L = 1.04 atm
P2 (CO2) = 2.15 atm x 4.87 L / 7.82 L = 1.34 atm
P (mixture) = 1.04 + 1.34 = 2.38 atm
Answer: 2.38 atm.